OpenStudy (anonymous):

Julie divided x^4 + 3x^3 -9x^2 +6 by a polynomial. Her answer was x^3 - 2x^2 + x - 5 with a remainder of 31. What polynomial did Julie divide by?

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

How would I find the divisor? Would I use the equation Dividend - Remainder / Quotient ?

OpenStudy (anonymous):

let the polynomial be ax+b then x^4 +3x^3-9x^2+6=(ax+b)(x^3-2x^2+x-5)+31 solve R.H.S and equate the terms with like powers.

OpenStudy (anonymous):

Do I expand (ax + b ) with (x^3-2x^2+x-5) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I get ax^4 + 2ax^3 + ax ^2 -5 ax + bx^3 - 2bx^2 +bx - 5b + 3. What do I do now?

OpenStudy (anonymous):

ax^4+(2a+b)x^3+(...........)x^2+(......)x+3+31=x^4+3x^3-9x^2+(.....)x+6 equate the terms with like powers and then find a and b.

OpenStudy (anonymous):

Thank you so much!